# General approach for partitioning and phase separation in macromolecular coexisting phases

### v1

@@ III.1 Bulk phase properties
Figure 5: System interacting as in Fig. 3. Here $\rho_{\rm r}=0.27$ and $\bar{\rho}_{\rm c}=0.09$ are fixed. In (a), phase response, and in (b) partitioning of the system components for varying $\bar{\rho}_{\rm a}$. In (c), phase response, and in (d) partitioning of the system components for varying $\bar{\rho}_{\rm b}$. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars, where their color corresponds to the value of $\bar{\rho}_{\rm a}$ (panel (a)) and $\bar{\rho}_{\rm b}$ (panel (c)). The green squares represent the value of $V_{\rm f}$, i.e. the volume ratio of phase-I, while the red squares represent $V_{\rm f}=0.5$. The differences in the coexisting densities of the colloids, $\Delta\rho_{\rm c}$, are shown in the insets of panel (b) and (d). Dashed line shown in panel (b) and (d) represents $N/N_{\rm 0}=0.5$.
Figure 6: System interacting as in Fig. 3 with the exception of $R_{ij}$ following Eq. (2). Here $\rho_{\rm r}=0.27$ and $\bar{\rho}_{\rm c}=0.09$ are fixed. In (a), phase response, and in (b) partitioning of the system components for varying $\bar{\rho}_{\rm a}$ under the constraint $\varphi_{\rm a}=\frac{4}{3}\pi R_{\rm aa}^{3}\bar{\rho}_{\rm a}=3.2$ and $\varphi_{\rm b}=\frac{4}{3}\pi R_{\rm bb}^{3}\bar{\rho}_{\rm b}=3.2$ with varying $R_{\rm aa}$. In (c), phase response, and in (d) partitioning of the system components for varying $\bar{\rho}_{\rm b}$ under the constraint $\varphi_{\rm a}=\frac{4}{3}\pi R_{\rm aa}^{3}\bar{\rho}_{\rm a}=3.2$ and $\varphi_{\rm b}=\frac{4}{3}\pi R_{\rm bb}^{3}\bar{\rho}_{\rm b}=3.2$ with varying $R_{\rm bb}$. Note that $\sigma_{\rm bb}=\sigma{R_{\rm bb}}/{R_{\rm r}}$, where $R_{\rm r}=\sqrt{2}\sigma$. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars, where their color corresponds to the value of $\bar{\rho}_{\rm a}$ (panel (a)) and $\bar{\rho}_{\rm b}$ (panel (c)). The green squares represent the value of $V_{\rm f}$, i.e. the volume ratio of phase-I, while the red squares represent $V_{\rm f}=0.5$. The differences in the coexisting densities of the colloids, $\Delta\rho_{\rm c}$, are shown in the insets of panel (b) and (d). Dashed line shown in panel (b) and (d) represents $N/N_{\rm 0}=0.5$.
@@
Figure 7: Here $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}=0.27$ and $\bar{\rho}_{\rm c}=0.09$ with $\beta\hat{V}=-5.6$. Coexistence densities (a) and partitioning (b) of system components for varying sizes of polymer a, where $R_{\rm bb}=\sqrt{2}\sigma=R_{\rm r}$. Note that $R_{\rm ab}$ follows from Eq. (2). Coexistence densities (c) and partitioning (d) of system components for varying sizes of polymer b, where $R_{\rm aa}=\sqrt{2}\sigma=R_{\rm r}$. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars, where their color corresponds to the value of $R_{\rm aa}$ and $R_{\rm bb}$ in panel (a) and (c), respectively. The green squares represent the value of $V_{\rm f}$, i.e. the volume ratio of phase-I, while the red squares represent $V_{\rm f}=0.5$. Full circle represents the critical point. Dashed line shown in panel (b) and (d) represents $N/N_{\rm 0}=0.5$.
@@
Finally, in order to explore the effect of asymmetry in polymer sizes, corresponding to asymmetry in depletion forces, we change the size of one polymer species while fixing the size of the other. Also here, we keep the average densities fixed as before ($\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}=0.27$ and $\bar{\rho}_{\rm c}=0.09$) and the interaction strength between colloids and polymer a at $\hat{V}_{\rm ac}=-5.6$. With $R_{\rm r}=\sqrt{2}\sigma$ being the reference polymer size, in Fig. 7(a),(b) we investigate the effect of $R_{\rm aa}$ on phase separation and partitioning, respectively, while keeping $R_{\rm bb}=R_{\rm r}$. Note that any change in polymer size implies a change in the value of $\sigma_{ii}$, which is linearly rescaled correspondingly, i.e. $\sigma_{ii}=\sigma R_{ii}/R_{\rm r}$, where $i=\rm a,b$. In panel (a) we see that the coexisting densities of the polymers are practically unaffected by the value of $R_{\rm aa}$, while the phase boundary, described by $V_{\rm f}$, shifts. Specifically, when $R_{\rm aa}>R_{\rm r}$, phase-I (polymer a-rich phase) increases in volume. Nevertheless, the partitioning of each species (shown in panel (b)), remain practically unaffected, with a slight variation for small values of $R_{\rm aa}$ in $N/N_{0}$ for polymer a. On the other hand, the difference in coexisting densities of the colloids decreases importantly as $R_{\rm aa}$ is increased. Similarly, in Fig. 7(c),(d) we study the same response yet as a function of $R_{\rm bb}$. Also in this case, as shown in panel (c), the values of the coexisting densities are barely affected, yet the phase boundary shifts such that the volume of phase-II increases as $R_{\rm bb}$ is increased. Also similar to the previous case, the partitioning of the different system components are practically unchanged (panel (d)). On the other, the inset of panel (d) shows that the difference of coexisting densities of the colloids correlates with the size of polymer b.
@@
Figure 8: Here $R_{\rm aa}=R_{\rm bb}=R_{\rm ab}=\sqrt{2}\sigma$, $\epsilon_{\rm aa}=\epsilon_{\rm bb}=2.0,\epsilon_{\rm ab}=2.5$, $\sigma_{\rm aa}=\sigma_{\rm bb}=\sigma$ and $\hat{V}_{\rm ac}=-5.6$. Additional constraints include $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}=0.27$, and reference density $\rho_{\rm r}=0.09$. In (a), phase response, and in (b) partitioning of the system components for varying $\bar{\rho}_{\rm c}$. In (c), phase response, and in (d) partitioning of the system components for varying $\bar{\rho}_{\rm c}$ with fix volume fraction $\varphi_{\rm c}=\frac{4}{3}\pi(\sigma_{\rm cc}/2)^{3}\bar{\rho}_{\rm c}=0.04712$, with $\bar{\rho}_{\rm c}=0.09$ ($\sigma_{\rm cc}$ varies accordingly). In (e), phase response, and in (f) partitioning of the different system components, for $\bar{\rho}_{\rm c}=0.09$ for varying colloids size, $\sigma_{\rm cc}$. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars (see colorbar). The green squares represent the value of $V_{\rm f}$, while the red squares represent $V_{\rm f}=0.5$. The differences in the coexisting densities of the colloids, $\Delta\rho_{\rm c}$ shown in the insets of panel (b), (d) and (f). Dashed lines shown in panel (b), (d) and (f) represents $N/N_{\rm 0}=0.5$.
@@ IV Conclusions and perspective
In this work, we proposed a general scheme for studying the bulk multiphase coexistence for a mixture composed by n polymer and m colloidal species with different affinities. This scheme can be applied to bimolecular condensates, where several polymeric fluid phases coexist with other dispersed macromolecules. In the scheme, we proposed an enhanced mean field approach within the classical density functional theory formalism Evans (1979). For the case of two demixing polymer species and one dispersed colloidal component affine to only one polymer type, our scheme has been benchmarked with particle-based simulations. We found both qualitative and quantitative results (see Figs. 2 and 9). This specific case is relevant in the context of water purification by means of aqueous two-phase systems Iqbal et al. (2016); Assis et al. (2021). We found that the partitioning of the colloidal species is highly sensitive to the strength of affinity towards the affine polymer species, as well as its density (the highest the density, the strongest the partitioning), together with the size ratio between polymers and colloids (smaller colloids partition more). On the other hand, we found marginal influence on the partitioning of colloids coming from variations in the density of the non-affine polymer species, and from the asymmetry in the two polymer species sizes.
@@ Acknowledgments
This work was supported by the Jane and Aatos Erkko Foundation under the grant No. 230052 (A.S.). Computational resources by CSC IT Center for Finland are also gratefully acknowledged. We are thankful to Prof. Andy Archer for fruitful discussions.

### v2

@@ III.1 Bulk phase properties
Figure 5: System interacting as in Fig. 3. Here $\rho_{\rm r}=0.27$ and $\bar{\rho}_{\rm c}=0.09$ are fixed. In (a), phase response, and in (b) partitioning of the system components for varying $\bar{\rho}_{\rm a}$. In (c), phase response, and in (d) partitioning of the system components for varying $\bar{\rho}_{\rm b}$. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars, where their color corresponds to the value of $\bar{\rho}_{\rm a}$ (panel (a)) and $\bar{\rho}_{\rm b}$ (panel (c)). The green squares represent the value of $V_{\rm f}$, i.e. the volume ratio of phase-I, while the red squares represent $V_{\rm f}=0.5$. The differences in the coexisting densities of the colloids, $\Delta\rho_{\rm c}$, are shown in the insets of panel (b) and (d). Dashed horizontal lines shown in panel (b) and (d) represent $N/N_{\rm 0}=0.5$.
Figure 6: System interacting as in Fig. 3 with the exception of $R_{ij}$ following Eq. (2). Here $\rho_{\rm r}=0.27$ and $\bar{\rho}_{\rm c}=0.09$ are fixed. In (a), phase response, and in (b) partitioning of the system components for varying $\bar{\rho}_{\rm a}$ under the constraint $\varphi_{\rm a}=\frac{4}{3}\pi R_{\rm aa}^{3}\bar{\rho}_{\rm a}=3.2$ and $\varphi_{\rm b}=\frac{4}{3}\pi R_{\rm bb}^{3}\bar{\rho}_{\rm b}=3.2$ with varying $R_{\rm aa}$. In (c), phase response, and in (d) partitioning of the system components for varying $\bar{\rho}_{\rm b}$ under the constraint $\varphi_{\rm a}=\frac{4}{3}\pi R_{\rm aa}^{3}\bar{\rho}_{\rm a}=3.2$ and $\varphi_{\rm b}=\frac{4}{3}\pi R_{\rm bb}^{3}\bar{\rho}_{\rm b}=3.2$ with varying $R_{\rm bb}$. Note that $\sigma_{\rm bb}=\sigma{R_{\rm bb}}/{R_{\rm r}}$, where $R_{\rm r}=\sqrt{2}\sigma$. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars, where their color corresponds to the value of $\bar{\rho}_{\rm a}$ (panel (a)) and $\bar{\rho}_{\rm b}$ (panel (c)). The green squares represent the value of $V_{\rm f}$, i.e. the volume ratio of phase-I, while the red squares represent $V_{\rm f}=0.5$. The differences in the coexisting densities of the colloids, $\Delta\rho_{\rm c}$, are shown in the insets of panel (b) and (d). Dashed horizontal lines shown in panel (b) and (d) represent $N/N_{\rm 0}=0.5$.
@@
Figure 7: Here $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}=0.27$ and $\bar{\rho}_{\rm c}=0.09$ with $\hat{V}_{\rm ac}=-5.6$. Coexistence densities (a) and partitioning (b) of system components for varying sizes of polymer a, where $R_{\rm bb}=\sqrt{2}\sigma=R_{\rm r}$. Note that $R_{\rm ab}$ follows from Eq. (2). Coexistence densities (c) and partitioning (d) of system components for varying sizes of polymer b, where $R_{\rm aa}=\sqrt{2}\sigma=R_{\rm r}$. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars, where their color corresponds to the value of $R_{\rm aa}$ and $R_{\rm bb}$ in panel (a) and (c), respectively. The green squares represent the value of $V_{\rm f}$, i.e. the volume ratio of phase-I, while the red squares represent $V_{\rm f}=0.5$. Full circle represents the critical point. Dashed horizontal lines shown in panel (b) and (d) represent $N/N_{\rm 0}=0.5$.
@@
Finally, in order to explore the effect of asymmetry in polymer sizes, corresponding to asymmetry in depletion forces, we change the size of one polymer species while fixing the size of the other. Such variation is relevant, as it has been recently shown experimentally that system containing mostly short polymers exhibit a narrower two-phase region, together with a reduced density differences Kamo et al. (2025). Also here, we keep the average densities fixed as before ($\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}=0.27$ and $\bar{\rho}_{\rm c}=0.09$) and the interaction strength between colloids and polymer a at $\hat{V}_{\rm ac}=-5.6$. With $R_{\rm r}=\sqrt{2}\sigma$ being the reference polymer size, in Fig. 7(a),(b) we investigate the effect of $R_{\rm aa}$ on phase separation and partitioning, respectively, while keeping $R_{\rm bb}=R_{\rm r}$. Note that any change in polymer size implies a change in the value of $\sigma_{ii}$, which is linearly rescaled correspondingly, i.e. $\sigma_{ii}=\sigma R_{ii}/R_{\rm r}$, where $i=\rm a,b$. In panel (a) we see that the coexisting densities of the polymers are practically unaffected by the value of $R_{\rm aa}$, while the phase boundary, described by $V_{\rm f}$, shifts. Specifically, when $R_{\rm aa}>R_{\rm r}$, phase-I (polymer a-rich phase) increases in volume. Nevertheless, the partitioning of each species (shown in panel (b)), remain practically unaffected, with a slight variation for small values of $R_{\rm aa}$ in $N/N_{0}$ for polymer a. However, this small change points towards a narrowing of the phase separation region of the polymer mixture, in line with Ref. 53. On the other hand, the difference in coexisting densities of the colloids decreases importantly as $R_{\rm aa}$ is increased. Similarly, in Fig. 7(c),(d) we study the same response yet as a function of $R_{\rm bb}$. Also in this case, as shown in panel (c), the values of the coexisting densities are barely affected, yet the phase boundary shifts such that the volume of phase-II increases as $R_{\rm bb}$ is increased. Also similar to the previous case, the partitioning of the different system components are practically unchanged (panel (d)). On the other, the inset of panel (d) shows that the difference of coexisting densities of the colloids correlates with the size of polymer b.
@@
Figure 8: Here $R_{\rm aa}=R_{\rm bb}=R_{\rm ab}=\sqrt{2}\sigma$, $\epsilon_{\rm aa}=\epsilon_{\rm bb}=2.0,\epsilon_{\rm ab}=2.5$, $\sigma_{\rm aa}=\sigma_{\rm bb}=\sigma$ and $\hat{V}_{\rm ac}=-5.6$. Additional constraints include $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}=0.27$, and reference density $\rho_{\rm r}=0.09$. In (a), phase response, and in (b) partitioning of the system components for varying $\bar{\rho}_{\rm c}$. In (c), phase response, and in (d) partitioning of the system components for varying $\bar{\rho}_{\rm c}$ with fix volume fraction $\varphi_{\rm c}=\frac{4}{3}\pi(\sigma_{\rm cc}/2)^{3}\bar{\rho}_{\rm c}=0.04712$, with $\bar{\rho}_{\rm c}=0.09$ ($\sigma_{\rm cc}$ varies accordingly). In (e), phase response, and in (f) partitioning of the different system components, for $\bar{\rho}_{\rm c}=0.09$ for varying colloids size, $\sigma_{\rm cc}$. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars (see colorbar). The green squares represent the value of $V_{\rm f}$, while the red squares represent $V_{\rm f}=0.5$. The differences in the coexisting densities of the colloids, $\Delta\rho_{\rm c}$ shown in the insets of panel (b), (d) and (f). Dashed horizontal lines shown in panel (b), (d) and (f) represent $N/N_{\rm 0}=0.5$.
@@ IV Conclusions and perspective
In this work, we proposed a general scheme for studying the bulk multiphase coexistence for a mixture composed by n polymer and m colloidal species with different affinities. This scheme can be applied to bimolecular condensates Holehouse and Alberti (2025), where several polymeric fluid phases coexist with other dispersed macromolecules. In the scheme, we proposed an enhanced mean field approach within the classical density functional theory formalism Evans (1979). For the case of two demixing polymer species and one dispersed colloidal component affine to only one polymer type, our scheme has been benchmarked with particle-based simulations. We found both qualitative and quantitative results (see Figs. 2 and 9). This specific case is relevant in the context of water purification by means of aqueous two-phase systems Iqbal et al. (2016); Assis et al. (2021). We found that the partitioning of the colloidal species is highly sensitive to the strength of affinity towards the affine polymer species, as well as its density (the higher the density, the stronger the partitioning), together with the size ratio between polymers and colloids (smaller colloids partition more). On the other hand, we found marginal influence on the partitioning of colloids coming from variations in the density of the non-affine polymer species, and from the asymmetry in the two polymer species sizes.
Notably, our findings resonate with previous studies on colloidal systems in binary solvent mixtures near their critical point Edison et al. (2015a, b), which also reported rich phase behavior and highlighted the role of parameters such as colloid size, colloid wettability and distance from the critical point, in controlling phase behavior. Although the physical systems are quite different (binary solvents in their case and immiscible polymers in water in ours), the parallels highlight the broader relevance of these mechanisms across several soft matter systems. Another interesting case that shares similarities with our work is the recent study of ouzo Archer et al. (2024), which consists of a ternary mixture of water, alcohol and trans-anethole oil. In this work, the experimental bulk phase diagram and interfacial properties have been benchmarked against a lattice DFT approach.
@@ Acknowledgments
This work was supported by the Jane and Aatos Erkko Foundation under the grant No. 230052 (A.S.). Computational resources by CSC IT Center for Finland are also gratefully acknowledged. The work was conducted within the #SUSMAT profiling measure at the University of Turku, Finland. We are also thankful to Prof. Andy Archer for fruitful discussions.